# Sysformer: Safeguarding Frozen Large Language Models with Adaptive System Prompts

### v1

@@ 6.4 How do input embeddings affect the performance of Sysformer?
Finally, we analyze the effect of changing the representations of user prompt embedding used to transform the system prompt embedding. The default implementation of Sysformer uses the LLM’s token embedding matrix to obtain useful user prompt embeddings to help learn the transformation. To understand the impact of these input representations, we use two state-of-the-art sentence embedding models: Linq  [^9] https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral and SFR  [^10] https://huggingface.co/Salesforce/SFR-Embedding-Mistral to embed the user prompts and pass the embeddings into the Sysformer architecture. Figure 5 compares the embeddings with the default token embedding matrix of each LLM in the JailbreakBench dataset. We find that the performance remains stable across different embedding architectures, showing the highest overall performance by using the LLM-specific token embedding matrix instead of a generic sentence embedding model. In particular, we note that the harm refusal rate in Phi-3.5-mini significantly reduces by using other embedding models, which highlights that the general-purpose embeddings may not be well-suited for these models but for trained models such as Llama-3.1-8b, these embeddings are applicable.
@@ 7 Conclusion
In this work, we have introduced Sysformer, a transformer-based mechanism that dynamically adapts the system prompt based on the user input to enhance the safety of frozen LLMs. Across five open LLMs and two safety benchmarks, we demonstrate that Sysformer improves robustness without retraining or filtering the user-provided prompts. These results challenge the prevailing notion of a fixed system prompt and highlight the potential of adaptive system prompts for safer LLM behavior. Our study is focused on small- to mid-scale models due to computational constraints, leaving the exploration of behavior at scale for future works and studying whether a universal Sysformer-like projector can be designed to work for any LLM. Another computational limitation arises due to a polynomial cost with the system prompt, which can be prohibitive considering other cache optimizations. Concurrent research has shown that safety alignment benefits greatly by increasing the response token depth (Qi et al., 2025), which can also be explored in the context of Sysformer. Another work has also discovered cone-shaped refusal directions instead of a single direction that can motivate more sophisticated training paradigms of our method (Wollschläger et al., 2025). Beyond safety, we believe that prompt-space adaptation can also benefit broader applications like retrieval-augmented generation, where Sysformer-like architectures can be used to adapt the retrieved context representation onto the user query, yielding more effective and query-aware responses.
@@
Appendix A Extended Related Work
Lightweight Model Adaptation. A common way to adapt models without extensive training is low-rank update of a few parameters (Hu et al., 2022). Other parameter-efficient fine-tuning approaches have also been proposed focusing on making minimal changes to the parameters (Liu et al., 2022; Ding et al., 2023). Here, we are interested in instead keeping all the learned parameters frozen.
Safety Benchmarks. Curation of high-quality harmful and safe prompts along with representative metrics is critical to understand and evaluate our progress in achieving safety. Thus, various resources and datasets have been developed for a systematic and comprehensive evaluation of LLM safety approaches (Chao et al., 2024; Souly et al., 2024; Mazeika et al., 2024, 2023; Wei et al., 2024). While performance on some of these benchmarks have been found to be confounded with other capabilities of scale (Ren et al., 2024), we use them to show gains in a large variety of fairly smaller LLMs.
System prompt leakage. Recently, there have been independent efforts to expose vulnerabilities in black-box large language models by leaking or finding their system prompts  [^11] https://tinyurl.com/5n6ms8mf   [^12] https://github.com/jujumilk3/leaked-system-prompts. These efforts have shown that upon knowing a system prompt, attackers can jailbreak the LLM easily into generating harmful content. The proposed Sysformer architecture addresses this issue by training an adaptive mechanism that generates a system prompt embedding conditioned on the user prompt, rather than relying on a fixed natural language system prompt. This is because these parameters are arguably harder to leak than a linguistic system prompt.
@@ Appendix B Algorithm
[…]
@@
Appendix C Prompts
C.1 Evaluation
Refusal check prompt
Prompt: "{prompt}"
Response: "{response}"
Final class:
C.2 System prompt and refusal response
Appendix D Additional Experimental Details
D.1 Implementation details
@@ D.2 Jailbreaks
  9. 9.
@@
Out of these, we use PAIR, PAP persuasion, distractors, Style injection, refusal suppression, and GCG universal attacks to augment our training set for training the “Ours+JB” model.
Table 2: Best configuration used for each method. F denotes False, T denotes True.
LLM JailbreakBench StrongReject
$w_{ref}$ $w_{compl}$ $w_{class}$ $w_{recon}$ selfsafe add $w_{ref}$ $w_{compl}$ $w_{class}$ $w_{recon}$ selfsafe add
@@
       SystemEmbedder 1 1 1 0 F F 1 0.2 1 1 T F
       Sysformer (ours) 1 1 1 1 F F 1 1 1 1 F F
@@
       SystemEmbedder 1 0.2 1 1 T F 1 1 1 0 F F
       Sysformer (ours) 1 0.5 1 1 F F 1 0.5 1 1 F F
@@
       SystemEmbedder 1 0.5 1 1 F F 1 0.2 1 1 F F
       Sysformer (ours) 1 0.5 1 0 F F 1 0.5 1 1 F F
@@
       SystemEmbedder 1 1 1 1 F F 1 0 1 1 F T
       Sysformer (ours) 1 0.2 1 0 F F 1 1 1 1 F F
@@
       SystemEmbedder 1 0.2 1 0 F F 1 0 1 0 F T
       Sysformer (ours) 1 0 1 0 F T 1 0 1 0 F T
@@
Appendix E Additional Experiments
@@
Best hyperparamters Table 2 shows the configuration of each hyperparamter to train each method that gives the best performance, as shown in Table 1. We find that the best performance is model and benchmark-dependent and $w_{compl}=0.2$ is often seen as the best performance with selfsafe not often used to find the optimal value.
Memory analysis. Table 3 shows the GPU memory used during training and test time for each method in the JailbreakBench dataset. We find that Sysformer uses more memory than SystemEmbedder for almost all LLMs during training but the additional memory is within $50$ GB while the memory during test time remains comparable. We argue that this gain in GPU memory is reasonable given the gains in performance.
Time analysis. Table 4 shows the time taken during inference of the whole JailbreakBench dataset and we find that Sysformer takes comparable times with SystemEmbedder in all cases. Note that train times cannot be compared across methods as the best configuration can be different, inculcating additional costs. However, we find that the gain in test time compared to the default system prompt is almost double. While this can be prohibitive at times, we argue that the gains in performance can justify this time. Furthermore, more optimizations can be done to further reduce this test time which has not been attempted in this work while our time complexity analysis also shows that the gains do not scale with the size of the model which addresses any scalability concerns.
@@
Jailbreaking attacks. We also provide the performance difference of Jailbreak-augmented dataset on the in and out of distribution jailbreaks separately. Here, in-distribution means the jailbreaking methods that were augmented during training and out-of-distribution denotes the others. We find no notable difference in the refusal rate for the two, showing great generalizability.
Strong Reject. Figure 8 provides compares different methods on StrongReject. We do not include “Ours+JB” here for brevity.
@@
Table 3: Total GPU memory allocated (in GiB) during training and testing the JailbreakBench.
@@
Table 4: Total time taken (s) for inference with different methods on the whole JailbreakBench dataset along with all the jailbreaks. Best train times can vary due to additional dataset in some.
LLM Method Total runtime (s)
zephyr-7b-beta DefaultSystem 814.7859
SystemEmbedder 2337.0526
Sysformer 2189.1768
Llama-2-7b-chat DefaultSystem 1037.8206
SystemEmbedder 2114.3058
Sysformer 2118.7483
Llama-3.1-8b DefaultSystem 835.2978
SystemEmbedder 2183.6973
Sysformer 2194.7296
Phi-3.5-mini DefaultSystem 686.0706
SystemEmbedder 1786.7062
Sysformer 1761.5489
Mistral-7B-v0.2 DefaultSystem 1069.1441
SystemEmbedder 2179.1167
Sysformer 2200.5760
@@
Figure 7: Comparison of Sysformer for total set of hyperparameters on JailbreakBench.
@@
Figure 9: Performance comparison of in and out of jailbreaking attacks on the JailbreakBench dataset.

### v2

@@ 6.5 How do input embeddings affect the performance of Sysformer?
Finally, we analyze the effect of changing the representations of the user prompt embedding used to transform the system prompt embedding. The default implementation of Sysformer uses the LLM’s token embedding matrix to obtain useful user prompt embeddings to help learn the transformation. To understand the impact of these input representations, we use two state-of-the-art sentence embedding models: Linq  [^12] https://huggingface.co/Linq-AI-Research/Linq-Embed-Mistral and SFR  [^13] https://huggingface.co/Salesforce/SFR-Embedding-Mistral to embed the user prompts and pass the embeddings into the Sysformer architecture. Figure 5 compares the embeddings with the default token embedding matrix of each LLM in the JailbreakBench dataset. We find that the performance remains stable across different embedding architectures, showing the highest overall performance by using the LLM-specific token embedding matrix instead of a generic sentence embedding model. In particular, we note that the harm refusal rate in Phi-3.5-mini significantly reduces by using other embedding models, which highlights that the general-purpose embeddings may not be well-suited for these models, but for trained models such as Llama-3.1-8b, these embeddings are applicable.
@@ 7 Conclusion
We introduce Sysformer, a transformer-based mechanism that dynamically adapts the system prompt to enhance the safety of frozen LLMs. Sysformer is found to boost the robustness without retraining or filtering, challenging the notion of a fixed system prompt and showing the potential of adaptive prompts for safer LLM behavior. Beyond safety, Sysformer can also inspire broader adaptive applications for other domains, such as retrieval-augmented generation, where adaptive projection aligns retrieved context with user queries. Concurrent work shows that deeper responses improve safety alignment (Qi et al., 2025), and that refusal directions are cone-shaped rather than singular (Wollschläger et al., 2025), both suggesting avenues for extending Sysformer.
Limitations. Our study is limited to small- and mid-scale models due to computational constraints, leaving scaling and universal plug-and-play projectors for future work. The method incurs polynomial costs with prompt length, which may hinder efficiency in certain cases and should inspire specific cache-optimized extensions. Finally, we note that Sysformer’s adaptivity could potentially introduce new vulnerabilities, as user prompts can directly influence the system prompt. However, such attacks are non-trivial to formulate in comparison with corresponding text-level ones since Sysformer operates on shallow embedding-level features rather than higher-level semantic reasoning. Regardless, investigating and mitigating such risks is an important direction for future work.
Declaration on LLM Usage
We use LLMs solely for revising the writing and framing of the text, and not in any other capacity.
Ethics statement
We use publicly-available benchmarks for harm-inducing prompts and adhere to their intended usage. Our contributions focus on mitigating these harms through trainable modular attachments to existing LLMs, and we do not identify additional ethical concerns. Nevertheless, we emphasize that our method should only be trained and applied in accordance with safe and ethically curated datasets.
@@
Appendix A Prompts
A.1 Evaluation
Refusal check prompt
Prompt: "{prompt}"
Response: "{response}"
Final class:
A.2 System prompt and refusal response
@@ Appendix B Algorithm
14:    $\mathcal{R}_{i}\leftarrow\text{``I am sorry I cannot help you."}$
@@
Appendix C Additional Experimental Details
C.1 Implementation details
@@ C.2 Jailbreaks
    Few-shot JSON: The model is given few-shot harmful question-answer pairs, and the user’s prompt is presented in JSON format, which the model is asked to continue generating.
  6. 6.
@@
  10. 10.
    PAP Persuasion: Instructs an attacker to persuade a victim model to respond using evidence-based persuasion.
  11. 11.
@@
Out of these, we use PAIR, PAP persuasion, distractors, Style injection, refusal suppression, and GCG universal attacks to augment our training set for training the “Ours+JB” model. For the full ood attacks in Table 3, refer  Souly et al. (2024).
Table 5: Best configuration used for each method in JailbreakBench. F denotes False, T denotes True.
$w_{ref}$ $w_{compl}$ $w_{class}$ $w_{recon}$ selfsafe add RR safe RR harm LG harm LG safe
@@
     SystemEmbedder 1 1 1 0 F F 0.0667 0.4000 0.9333 0.2000
     Sysformer (ours) 1 1 1 1 F F 0.1667 0.9333 0.8667 0.8000
@@
     SystemEmbedder 1 0.2 1 1 T F 0.5667 1.0000 0.9333 1.0000
     Sysformer (ours) 1 0.5 1 1 F F 0.0667 0.9000 0.9000 0.8667
@@
     SystemEmbedder 1 0.5 1 1 F F 0.3000 1.0000 1.0000 1.0000
     Sysformer (ours) 1 0.5 1 0 F F 0.0333 0.9667 0.8333 0.9667
@@
     SystemEmbedder 1 1 1 1 F F 0.0333 0.1667 0.6667 0.0667
     Sysformer (ours) 1 0.2 1 0 F F 0.2000 0.9000 0.8667 1.0000
@@
     SystemEmbedder 1 0.2 1 0 F F 0.1333 0.8667 0.9333 0.9333
     Sysformer (ours) 1 0 1 0 F T 0.1000 1.0000 0.9333 1.0000
Table 6: Best configuration used for each method in StrongReject. F denotes False, T denotes True.
$w_{ref}$ $w_{compl}$ $w_{class}$ $w_{recon}$ selfsafe add RR safe RR harm LG harm LG safe
zephyr-7b-beta
     SystemEmbedder 1 0.2 1 1 T F 0.0667 0.3404 0.9333 0.3191
     Sysformer (ours) 1 1 1 1 F F 0.1333 0.7553 0.8667 0.6170
Llama-2-7b-chat
     SystemEmbedder 1 1 1 0 F F 0.0667 0.4000 0.9333 0.2000
     Sysformer (ours) 1 0.5 1 1 F F 0.0333 0.8085 0.9333 0.8085
Llama-3.1-8b
     SystemEmbedder 1 0.2 1 1 F F 0.3000 1.0000 1.0000 1.0000
     Sysformer (ours) 1 0.5 1 1 F F 0.0333 1.0000 0.9000 1.0000
Phi-3.5-mini
     SystemEmbedder 1 0 1 1 F T 0.0667 0.2660 0.6667 0.0319
     Sysformer (ours) 1 1 1 1 F F 0.0667 0.5851 0.9000 0.8617
Mistral-7B-v0.2
     SystemEmbedder 1 0 1 0 F T 0.1333 0.9362 0.9333 0.9574
     Sysformer (ours) 1 0 1 0 F T 0.1000 1.0000 0.9333 0.9681
@@
Appendix D Additional Experiments
@@
Best hyperparameters Tables 5 and  6 show the configuration of each hyperparameter to train each method that gives the best performance, as shown in Table 1. We find that the best performance is model and benchmark-dependent, and $w_{compl}=0.2$ is often seen as the best performance with selfsafe not often used to find the optimal value.
Memory analysis. Table 8 shows the GPU memory used during training and test time for each method in the JailbreakBench dataset. We find that Sysformer uses more memory than SystemEmbedder for almost all LLMs during training, but the additional memory is within $50$ GB while the memory during test time remains comparable. We argue that this gain in GPU memory is reasonable given the gains in performance.
@@
Jailbreaking attacks. We also provide the performance difference of the Jailbreak-augmented dataset on the in and out of distribution jailbreaks separately in Figure 11. Here, in-distribution means the jailbreaking methods that were augmented during training, and out-of-distribution denotes the others. We find no notable difference in the refusal rate for the two, showing great generalizability.
Strong Reject. Figure 9 provides a comparison of different methods on StrongReject. We do not include “Ours+JB” here for brevity.
@@
Performance of suffix optimization. Here, we consider a version of the contemporaneous work (Li et al., 2025) as an additional baseline. In particular, we train a prompt suffix (with 5 token length) in the embedding space (similar to how SystemEmbedder trains the system prompt) using our set of loss functions for a fair comparison. Note that the main difference is that while SOP uses coordinate descent and optimizes in the token space, we use gradient descent and optimize in the embedding space. Despite these differences, we believe that our considered baseline captures some key aspects of SOP, specifically the placement of its embedding and adaptive nature due to the causal mask.
Table 7 shows the results of this method for the JailbreakBench dataset, and we find that it performs worse than the methods considered in our work (with a $\Delta$ $R$ $\leq 0.6$). Specifically, while it reduces the refusal rate on safe prompts, it cannot simultaneously increase the refusal rate on harmful prompts. Sysformer thus outperforms the refusal gap by a large margin.
LLM RR safe RR harm $\Delta$ $R$
zephyr-7b 0.0000 0.5000 0.5000
Llama-2-7b 0.1667 0.7333 0.5667
Llama-3.1-8b 0.0000 0.6000 0.6000
phi-3.5-mini 0.0667 0.5667 0.5000
Mistral-7B-v0.2 0.0000 0.6000 0.6000
Table 7: Performance of suffix-optimization baseline on JailbreakBench.
Training loss curves. Figure 7 shows the training curves of different loss functions as the model is trained on a combination of these. Other models also show similar trends, highlighting the training stability and importance of each loss.
((a))
((b))
((c))
((d))
Figure 7: Training curves of Llama-3.1-8B with $w_{ref}=1,w_{compl}=0.5,w_{class}=1,w_{recon}=1$.
Table 8: Total GPU memory allocated (in MB) during training and testing the JailbreakBench.
@@
Figure 8: Comparison of Sysformer for the total set of hyperparameters on JailbreakBench.
@@
Figure 10: Comparison of different methods on JailbreakBench by using a different pad token.
((a))
((b))
((c))
((d))
((e))
Figure 11: Performance comparison on in and out of distribution attacks on the JailbreakBench dataset for the augmented Ours+JB method.